OpenStudy (anonymous):

\[\LARGE \log(\log x)+\log(\log x^5-4)=0\] A) x=1 B) x=-1 C) x=-2 D) x=0

OpenStudy (experimentx):

let log x = u

OpenStudy (anonymous):

... let me see what I can do. :)

OpenStudy (anonymous):

but then ... \[\LARGE \log(u)+\log(5u-4)=0\] \[\LARGE \log(5u^2-4u)=0\] now according to: \[\LARGE \log_ab=c \quad \Longrightarrow b=a^c\] I thought about: \[\LARGE 5u^2-4u=10^0\] \[\LARGE 5u^2-4u-1=0\] \[\LARGE u_{1/2}=\frac{4\pm\sqrt{16-20\cdot (-1)}}{10}\] \[\LARGE u_{1/2}=\frac{4\pm6}{10}\] \[\LARGE u_1=1 \quad \quad ,\quad \quad u_2=-\frac25\] so... \[\LARGE \log x=1 \Longrightarrow x=10\] and \[\LARGE \log x=-\frac25 \Longrightarrow x=\frac{1}{\sqrt[5]{100}}\] ... O_O

OpenStudy (anonymous):

sorry.. it's -1/5 \[\LARGE x=\frac{1}{\sqrt[5]{10}}\]

OpenStudy (anonymous):

@satellite73 since you love logs :P , please can you have a look what I'm doing wrong :(

OpenStudy (anonymous):

why it says "e" to me ! ?:O http://www.wolframalpha.com/input/?i=+log%28log+x%29%2Blog%28log+x%5E5-4%29%3D0

OpenStudy (experimentx):

use this log[10, x]

OpenStudy (anonymous):

but it says there that is in base 10 !... it doesn't matter, I don't know if my PC went crazy, but your link doesn't show a solution or something. O_O , :( ..@experimentX :)

OpenStudy (experimentx):

your method is right .. I think so.

OpenStudy (anonymous):

but my options... :(

OpenStudy (experimentx):

put this on wolfram alpha. solve log[10, log [10, x]] + log[10, (log[10, x^5] - 4)]

OpenStudy (anonymous):

x=10 ? :O

OpenStudy (mertsj):

No.

OpenStudy (anonymous):

then how... can you provide a better solution, that's what I'm looking for.

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Wolf says 2.80688

OpenStudy (mertsj):

So based on the posted problem and the posted answers, I'd say there is an error in the posting.

OpenStudy (mertsj):

Make sure when you type it into wolf that you use the base 10 option.

OpenStudy (anonymous):

ok hold on a second...

OpenStudy (anonymous):

Question Nr. 7

OpenStudy (mertsj):

Well, just try replacing x with each of those answers. Which one works?

OpenStudy (mertsj):

Every single one of them results in a negative argument in log(x^5-4)

OpenStudy (anonymous):

I don't know :( I'm pretty lost \[\LARGE \log_{10}(0)=?\] since: \[\LARGE \log_ab \quad\quad \quad ,\quad \quad b>0\] so I guess it can't be option D :O

OpenStudy (mertsj):

10^0=1 but 10 ^x cannot be 0

OpenStudy (mertsj):

log 0 is undefined.

OpenStudy (mertsj):

I see this is in a foreign language. Perhaps the notation means something different. Why do you need to do this problem anyway?

OpenStudy (anonymous):

just to exercise , nothing special ^_^ ... It's ok, I'll ask my professor probably there's a mistake in the book. ;) Thanks for the help @Mertsj @experimentX

OpenStudy (experimentx):

yw

OpenStudy (mertsj):

yw When you find out, you might enlighten us as well.

OpenStudy (anonymous):

lol , ok. Although I'm sure there must be a mistake ;) . Have a nice day .